### v1

@@ 2.2 Detection probability and objectness score
Figure 3: Illustration of the output vector of a single detection unit. The elements are colored by type. The corresponding activation functions are indicated. For multiple detection units per grid cell, this vector structure is repeated on the same axis (Sect. 2.5).
To obtain a working object detector, we not only have to predict bounding boxes but also to evaluate the chances that they indeed contain an object. For this, we add a self-assessed detection probability prediction $P$ to each detection unit, which is constrained during training. This term uses a sigmoid activation and adds a sum-of-square error contribution to the loss. Due to our grid structure, we have one possible box per grid cell. In a context with only a few target boxes in the image, most of the grid cells map irrelevant background regions. During training, we identify the predicted boxes that best represent each target box and attribute them a target probability of $\hat{P}=1$. For all the remaining empty predicted boxes, we associate a target probability of $\hat{P}=0$. To compensate for the probable imbalance between the number of matching and empty predictions, we must define a usually small $\lambda_{void}$ factor to apply to the loss term representing empty predicted boxes. This helps balance the contribution of the two terms. The resulting loss term can be written as
$\mathcal{L}_{prob}=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{match}(P_{i}-1)^{2}+\mathds{1}_{i}^{void}\lambda_{void}(P_{i}-0)^{2},$ (8)
where the sum over $i$ represents all the grid cells, $\mathds{1}_{i}^{match}$ is a mask to identify the predicted boxes that match a target box, and $\mathds{1}_{i}^{void}$ a mask to identify the empty predicted boxes. Due to the stochasticity of the training process, it should result in a continuous probability distribution. At prediction time, the probability is used to identify the grid cells that should contain an object.
This probability definition contains no information about how well the predicted box represents the object. To account for this, we must define a metric that measures the proximity and resemblance between two bounding boxes. The classical metric for object detection is the intersection over union (IoU, Everingham et al., 2010; Lin et al., 2014). It is defined as the surface area of the intersection between two boxes, A and B, divided by the surface area of their union, which is expressed as
@@
This quantity takes values between 0 and 1 depending on the amount of overlap. The IoU can then be used to select the best prediction for a given target and quantify the quality of the prediction, but more generally, it can be used to compare two boxes of any kind. This classical IoU is the most commonly used in computer vision, but it presents some weaknesses for astronomical applications. We present a few alternative matching metrics better suited to our application case in Appendix A.4. Because several hyper-parameters of our method depend on this choice of metric, we will use a generic ${\rm fIoU}$ term that can be replaced by the selected matching metric in all the following equations. The default choice for our detector is the distance-IoU (DIoU Zheng et al., 2019), as it includes information about the distance between the center of the two boxes to compare. For all cases where it matters, the selected metric function is linearly rescaled in the 0 to 1 range if it was not already the case.
@@
From this, we add a self-assessed score $O$ called ”objectness” to each predicted box, which is also constrained during training. The objectness is defined as the combination of an object presence probability $P_{obj}$ and the ${\rm fIoU}$ between the predicted box and the target box, expressed as
$O=P_{obj}\times{\rm fIoU}.$ (10)
@@
As for the probability, this new term uses a sigmoid activation and adds a sum-of-square error contribution to the loss. The objectness is constrained like the probability by considering that $\hat{P}_{obj}=1$ for prediction-target matches, while $\hat{P}_{obj}=0$ for empty predicted boxes. The difference is the target objectness for the prediction that matches that is defined as $\hat{O}={\rm fIoU}$, following Eq.10, using the ${\rm fIoU}$ of the identified prediction-target couple. The resulting loss term can be written as
@@
$\mathcal{L}_{obj}=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{match}(O_{i}-{\rm fIoU})^{2}+\mathds{1}_{i}^{void}\lambda_{void}(O_{i}-0)^{2},$ (11)
@@
using the same notations as for Eq. 8. We stress that ${\rm fIoU}$ is used as a scalar in this equation. The derivative of the corresponding matching function is not computed for gradient propagation, so $\mathcal{L}_{obj}$ does not contribute to updating the position and size of the prediction. After training, we should obtain a continuous objectness distribution representing a global detection score that includes a self-assessment of the predicted box quality. We note that the classical YOLO formalism only predicts objectness, while YOLO-CIANNA predicts both probability and objectness. They can then be used independently or in association to construct advanced prediction filtering conditions (Sect. 2.8).
@@
With this formalism, we formulate only two statuses for a predicted box, either a match or empty, while in practice, multiple predicted boxes can try to represent the same target simultaneously. This is common if the target box center is positioned at the edge of a grid cell or if the boxes are large. This will be even more common with multiple detections per grid cell (Sect. 2.5). In such a case, only the best-predicted box will be considered a match. The remaining plausible detections are called good-but-not-best (GBNB) predictions. The previous formalism would result in a loss that lowers the objectness of these GBNB predictions, actively forcing relevant features to fade. To prevent this, we define a representation quality threshold ${\rm fIoU}\geq L^{{\rm fIoU}}_{gbnb}$ above which the corresponding boxes are excluded from both $\mathds{1}_{i}^{match}$ and $\mathds{1}_{i}^{void}$ masks. In summary, there are three types of contribution to the loss: i) the best detection for each target updates its box position and size while increasing its probability and objectness, ii) the background boxes lower their probability and objectness, and iii) the GBNB boxes are ignored.
@@ 2.3 Classification
The detected box can be enriched with a classification capability. With the classical YOLO formalism, it can be done by adding $N_{c}$ components, corresponding to all the possible classes, to the output vector of the detected boxes. The activation of these components can either be i) a sigmoid for all classes using a sum-of-square error, which allows multi-labeling, or ii) a soft-max activation, which corresponds to exponentiating all the outputs and normalizing them so their sum is equal to 1, with a cross-entropy error. These two options are available in our method. In both cases, only the best detection for each target box updates its classes by comparing the target class vector with the predicted one. There is no contribution to the class loss from both GBNB and background predictions. The resulting loss term for a soft-max activation with a cross-entropy error can be written as
$\mathcal{L}_{class}=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{match}\sum_{k}^{N_{C}}\left(-\hat{C}_{i}^{k}\log(C_{i}^{k})\right),$ (12)
where the sum over $k$ represents all the classes for a given predicted box, and $C_{i}^{k})$ is the corresponding class output for the $k$-th class of the predicted box $i$. We note that classification was not used for the SDC1 as discussed in Sect. 3.1, but it is used for benchmarks on computer vision datasets in Appendix A.8.
@@ 2.4 Additional parameters prediction
For astrophysical applications, we usually need to predict source properties like the flux or some geometric properties not described by a bounding box formalism. For this, we propose to add $N_{p}$ components to the output vector of the detected boxes, corresponding to all the additional parameters to predict. The activation of these components is linear with a sum-of-square error contribution to the loss. The respective contribution of these parameters to the loss can be scaled with a set of $\gamma^{p}$ factors. The resulting loss term can be written as
@@
$\mathcal{L}_{param}=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{match}\sum_{k=0}^{N_{p}}\gamma^{k}\left(p^{k}_{i}-\hat{p}^{k}_{i}\right)^{2},$ (13)
@@
where the sum over $k$ represents all the independent parameters for a given predicted box, and $p_{i}^{k}$ is the corresponding parameter output for the $k$-th parameter of the predicted box $i$. We emphasize that it is a strong added value of our YOLO-CIANNA method, allowing it to predict an arbitrary number of added values per detection for any application while preserving the one-stage formalism specific to regression-based object detectors.
2.5 Multiple boxes per grid-cell
@@
With the present definition, the detector output would have a shape of $g_{w},g_{h},(6+N_{c}+N_{p})\right>$, where $g_{w}$ and $g_{h}$ are the grid dimensions, the six static parameters are the box coordinates, probability, and objectness ($x,y,w,h,P,O$), $N_{c}$ is the number of classes, and $N_{p}$ is the number of additional parameters. While the geometric and detection score parameters are always present, both $N_{c}$ and $N_{p}$ are problem-dependent and user-defined. The typical vector for each grid cell with highlighted sub-parts and the corresponding activation functions is illustrated in Fig. 3.
@@
Application cases for which only one object would have to be detected per grid element are uncommon, and high grid resolutions are computationally expensive (Appendix A.2). To overcome this, the classical YOLO approach expands the output vector at each grid cell to contain multiple boxes by stacking their independent vector as a longer 1D vector. The new output shape is then $g_{w},g_{h},N_{b}{\times}(6+N_{c}+N_{p})\right>$, with $N_{b}$ the number of independent boxes predicted by each grid cell. We define an individual size-prior ($p_{w},p_{h}$) for each possible box in a given grid cell, which impacts the size scaling in Eqs. 4 and 5. This definition helps to distribute objects over the available boxes on a given grid cell based on box sizes and shapes. The prior list is the same for each grid cell due to the fully convolutional network backbone required for our method (Appendix A.1). In the latter, we refer to these independent predictive elements as ”detection units”. For example, a network set capable of predicting up to 6 independent boxes per grid cell is composed of 6 detection units. Some detection units can have the same size prior, but they still represent independent predictions.
@@
Because scales or shapes are unlikely to be evenly represented in the training sample, the detection units should adapt their positive-to-negative detection ratio to rebalance their probability and objectness distribution (Sect. 2.2). For this, a $\lambda_{void}^{b}$ factor is defined for each detection unit. These factors must be adjusted so the objections and probability responses are sufficiently similar for all detection units to be compared during prediction filtering (Sect. 2.8).
@@ 2.6 Target-Prediction association function
With multiple box predictions per grid cell, it is necessary to establish a clear process to decide which detection unit should be associated with each target box. Inside our YOLO-CIANNA method, we introduce a new prediction-aware association process. Our approach is very different than the one used in the classical YOLO formalism, which only uses the size priors to define the theoretical best matches. Our method proposes solutions to identified limitations of the classical YOLO for astronomical applications, especially regarding images with a high density of small objects. An in-depth comparison of our approach with the classical YOLO is presented in Appendix A.3.
@@
The main objective of our association process is to find the best target-prediction pairs regarding a specific ${\rm fIoU}$ matching metric. We start by setting $\mathds{1}^{void}=1$ for all detection units. Then, we identify all predicted boxes that are a good enough representation of at least one target regarding our ${\rm fIoU}\geq L^{{\rm fIoU}}_{gbnb}$ threshold. This comparison is made for all detection units and all targets regardless of their center position. All objects that respect this criterion are removed from $\mathds{1}^{void}$. The rest of the association algorithm aims to find the ”best match” for each target box through an iterative process. First, matching scores for all possible target-prediction pairs in a grid cell are stored in a scoring matrix. Then, the best current score in the matrix is used to define a new target-prediction pair that is added to the $\mathds{1}^{match}$ mask and removed from $\mathds{1}^{void}$ if it was not already the case. The full row and column corresponding to the target and detection unit of the best match are masked in the score matrix. This search process is repeated until the score matrix is empty or fully masked. As presented in Sect. 2.2, the best match associations contribute to all sub-parts of the loss, while the detection units that remained in $\mathds{1}^{void}$ contribute only to the probability and objectness following Eq. 11. The remaining GBNB detection units not being part of either $\mathds{1}^{match}$ or $\mathds{1}^{void}$ do not contribute to the loss.
@@
To account for edge cases, size imbalance, or training difficulty, we added several specific refinements inside our custom YOLO-CIANNA association process, which we detail in an appendix section dedicated to advanced subtleties of our method (Appendix A.6). The global association function algorithm is presented in Fig. 4 in a way that separates the simplified association process and the advanced association with refinements.
Figure 4: Summary of the target-prediction association algorithm of YOLO-CIANNA. All the steps are done independently by each grid cell. All the elements are executed in order from top to bottom, with the left column representing the general case for the association. All the refinements corresponding to the B.2 block are described in Sect. A.6.

### v2

@@ 2.2 Detection probability and objectness score
Only predicting bounding boxes is insufficient to obtain an object detector. We also need to evaluate the chances that they contain an object. For this, we add a self-assessed detection probability prediction $P$ to each grid cell, which is constrained during training. This term uses a sigmoid activation and adds a sum-of-square error contribution to the loss. Due to our grid structure, the detector always outputs box properties for every grid cell. In a context where only a few target boxes are present in the image, most grid cells only map irrelevant background regions. During training, we identify the predicted boxes that best represent each target box and attribute them a target probability of $\hat{P}^{\,\textrm{match}}=1$. For all the remaining empty predicted boxes, we attribute them a target probability of $\hat{P}^{\,\textrm{void}}=0$. To compensate for the likely imbalance between the number of matching and empty predictions, we must define a $\lambda_{\textrm{void}}$ factor to apply to the loss term corresponding to the empty case. The resulting loss term can be written as
$\begin{split}\mathcal{L}_{\textrm{prob}}&=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\left(P_{i}-\hat{P}_{i}^{\,\textrm{match}}\right)^{2}+\mathds{1}_{i}^{\textrm{void}}\lambda_{\textrm{void}}\left(P_{i}-\hat{P}_{i}^{\,\textrm{void}}\right)^{2}\\ &=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\left(P_{i}-1\right)^{2}+\mathds{1}_{i}^{\textrm{void}}\lambda_{\textrm{void}}\left(P_{i}-0\right)^{2},\end{split}$ (8)
where the sum over $i$ represents all the grid cells, $\mathds{1}_{i}^{\textrm{match}}$ is a mask to identify the predicted boxes that match a target box, and $\mathds{1}_{i}^{\textrm{void}}$ a mask to identify the empty predicted boxes. Through the stochasticity of the training process, the network will learn to predict a continuous probability score that reflects the detection confidence. At prediction time, it is used to identify the grid cells that should contain an object.
One limit of this probability definition is that it contains no information about the quality of the predicted box. For this, we must define a metric that measures the proximity and resemblance between two bounding boxes. The classical object detection metric for this is the intersection over union (IoU, Everingham et al., 2010; Lin et al., 2014). It is defined as the surface area of the intersection between two boxes, A and B, divided by the surface area of their union, which is expressed as
@@
This quantity takes values between 0 and 1 depending on the amount of overlap. This classical IoU is the most commonly used in computer vision, but it presents some weaknesses for astronomical applications. We present a few alternative matching metrics better suited to our application case in Appendix A.4. Since several hyperparameters of our method depend on this choice of metric, we will use a generic ${\rm fIoU}$ term that can be replaced by the selected matching metric in all the following equations. The default choice for our detector is the distance-IoU (DIoU, Zheng et al., 2020), as it includes information about the distance between the center of the two boxes to compare. For the unicity of the equations, the selected match metric function is always considered linearly rescaled in the 0 to 1 range.
@@
From this, we add a self-assessed score called objectness, $O,$ to each predicted box, which is also constrained during training. The objectness is defined as the combination of an object presence probability $P$ and the ${\rm fIoU}$ between the predicted box and the target box, expressed as
$O=P\times{\rm fIoU}.$ (10)
@@
This term also uses a sigmoid activation and adds a sum-of-square error contribution to the loss. During training, the objectness is constrained like the probability by considering that $\hat{P}^{\,\textrm{match}}=1$ for prediction-target matches, while $\hat{P}^{\,\textrm{void}}=0$ for empty predictions. Therefore, following Eq. 10, the target objectness for prediction-target matches is $\hat{O}^{\textrm{match}}={\rm fIoU}$, using the ${\rm fIoU}$ between the target and predicted boxes. For predictions with no associated matches, the target objectness is $\hat{O}^{\textrm{void}}=0$. The resulting loss term can be written as
@@
$\begin{split}\mathcal{L}_{\textrm{obj}}&=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\left(O_{i}-\hat{O}_{i}^{\,\textrm{match}}\right)^{2}+\mathds{1}_{i}^{\textrm{void}}\lambda_{\textrm{void}}\left(O_{i}-\hat{O}_{i}^{\,\textrm{void}}\right)^{2}\\ &=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\left(O_{i}-{\rm fIoU}\right)^{2}+\mathds{1}_{i}^{\textrm{void}}\lambda_{\textrm{void}}\left(O_{i}-0\right)^{2},\end{split}$ (11)
@@
using the same notations as Eq. 8. We stress that ${\rm fIoU}$ is used as a scalar in this equation. Therefore, the derivative of the corresponding matching function is not computed for gradient propagation, so $\mathcal{L}_{\textrm{obj}}$ does not contribute to updating the position and the size. After training, we obtain a continuous objectness prediction representing a global detection score that accounts for the predicted geometrical box quality. Probability and objectness can be used independently or in association to construct advanced prediction filtering conditions (Sect. 2.8).
@@
With this formalism, we formulate only two states for a predicted box, either a match or empty. In practice, multiple predicted boxes can try to represent the same target simultaneously. This is common if the target box center is positioned at the edge of a grid cell or if the boxes are large. This will be even more common with multiple detections per grid cell (Sect. 2.5). In such a case, only the best-predicted box will be considered a match. The remaining plausible detections are called good-but-not-best (GBNB) predictions. The previous formalism would result in a loss that lowers the objectness of these GBNB predictions, actively forcing relevant features to fade. To prevent this, we define a representation quality threshold ${\rm fIoU}\geq L^{{\rm fIoU}}_{\textrm{GBNB}}$ above which the corresponding boxes are excluded from both $\mathds{1}_{i}^{\textrm{match}}$ and $\mathds{1}_{i}^{\textrm{void}}$ masks. In summary, there are three types of contribution to the loss: i) the best detection for each target updates its box position and size while increasing its probability and objectness, ii) the background boxes lower their probability and objectness, and iii) the GBNB boxes are ignored.
@@ 2.3 Classification
The detected box can be enriched with a classification capability. This can be done by adding $N_{c}$ components, corresponding to all the possible classes, to the output vector of the detected boxes (Fig. 3). The activation of these components can either be i) a sigmoid for all classes using a sum-of-square error, which allows multi-labeling, or ii) a soft-max activation, which corresponds to exponentiating all the outputs and normalizing them so their sum is equal to 1, with a cross-entropy error. These two options are available in our method. In both cases, only the best detection for each target box updates its classes by comparing the target class vector with the predicted one. There is no contribution to the class loss from either GBNB or background predictions. The resulting loss term for a soft-max activation with a cross-entropy error can be written as
$\mathcal{L}_{\textrm{class}}=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\sum_{k}^{N_{C}}\left(-\hat{C}_{i}^{k}\log(C_{i}^{k})\right),$ (12)
where the sum over $k$ represents all the classes for a given predicted box, and $C_{i}^{k})$ is the corresponding class output for the $k$-th class of the predicted box $i$.
@@ 2.4 Additional parameters prediction
For astrophysical applications, we often need to predict the characteristics of the sources, such as the flux or some geometric properties not described by a bounding box formalism. For this, we propose to add $N_{p}$ components to the output vector of the detected boxes, corresponding to all the additional parameters to predict. The activation of these components is linear with a sum-of-square error contribution to the loss. The respective contribution of these parameters to the loss can be scaled with a set of $\gamma^{p}$ factors. The resulting loss term can be written as
@@
$\mathcal{L}_{\textrm{param}}=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\sum_{k=0}^{N_{p}}\gamma^{k}\left(p^{k}_{i}-\hat{p}^{k}_{i}\right)^{2},$ (13)
@@
where the sum over $k$ represents all the independent parameters for a given predicted box, and $p_{i}^{k}$ is the corresponding parameter output for the $k$-th parameter of the predicted box $i$. This is a strong added value of our YOLO-CIANNA method compared to other approaches, allowing it to predict an arbitrary number of additional properties per detection for any application while preserving the one-stage formalism specific to regression-based object detectors.
2.5 Multiple boxes per grid-cell and detection unit definition
Figure 3: Illustration of the output vector of a single detection unit. The elements are colored according to the corresponding loss subpart. For multiple detection units per grid cell, this vector structure is repeated on the same axis (Sect. 2.5).
@@ 2.5 Multiple boxes per grid-cell and detection unit definition
With the present definition, the detector output would have a shape of $g_{w},g_{h},(6+N_{c}+N_{p})\right>$, where $g_{w}$ and $g_{h}$ are the grid dimensions, the six static parameters are the box coordinates, probability, and objectness ($x,y,w,h,P,O$), $N_{c}$ is the number of classes, and $N_{p}$ is the number of additional parameters (Fig. 3). While the geometric and detection score outputs are always predicted, both $N_{c}$ and $N_{p}$ are problem-dependent and user-defined.
@@
In theory, the network reduction factor could be adjusted to the average object density of the application to have no more than one target object in each grid cell. However, for many use cases, it would result in an output grid resolution close to the input resolution, which is very computationally intensive (Appendix A.2). To overcome this, we can expand the output vector at each grid cell to contain multiple boxes by stacking their independent vector as a longer 1D vector. The new output shape is then $g_{w},g_{h},N_{b}{\times}(6+N_{c}+N_{p})\right>$, with $N_{b}$ the number of independent boxes predicted at each grid cell, which is a hyperparameter of the detector. For each possible box in a grid cell, we define an individual size-prior ($p_{w},p_{h}$), which impacts the size scaling in Eqs. 4 and 5. This definition helps to distribute objects over the available boxes on a given grid cell based on box sizes and shapes. The prior list is the same for all grid cells as they only represent different positions at which the same detector is applied through the fully convolutional architecture (Appendix A.1). The size priors are user-defined hyperparameters but are often automatically defined through clustering of the size distribution of targets from the training sample. In the latter, we refer to the independent predictive elements for a single grid cell as detection units. For example, a detector set capable of predicting up to six independent boxes per grid cell comprises six detection units. Some detection units can have the same size prior, but they still represent independent predictions.
@@
Because scales or shapes are unlikely to be evenly represented in the training sample, the detection units should adapt their positive-to-negative detection ratio to rebalance their probability and objectness distribution (Sect. 2.2). For this, a $\lambda_{\textrm{void}}^{b}$ factor is defined for each detection unit. These factors must be adjusted so the objectness and probability responses are similar enough for all detection units to be comparable during prediction filtering (Sect. 2.8).
@@ 2.6 Target-prediction association function
With multiple box predictions per grid cell, deciding which detection unit should be associated with each target box is critical. With our YOLO-CIANNA method, we introduce a prediction-aware association process. Our approach differs significantly from the one used in the classical YOLO formalism, which only uses size priors to define theoretical best matches. Our method is expected to be more robust for images with a high density of small objects, which is typical for astronomical data products. An in-depth discussion about the motivation behind our implementation and a comparison with the classical YOLO association process are both presented in Appendix A.3.
@@
The main objective of our association process is to find the best target-prediction pairs regarding a specific ${\rm fIoU}$ matching metric. We start by setting $\mathds{1}^{\textrm{void}}=1$ for all detection units. Then, we identify all predicted boxes that are a good enough representation of at least one target regarding our ${\rm fIoU}\geq L^{{\rm fIoU}}_{\textrm{GBNB}}$ threshold. This comparison is made for all detection units and all targets regardless of their center position. All objects that respect this criterion are removed from $\mathds{1}^{\textrm{void}}$. The rest of the association algorithm aims to find the best match for each target box through an iterative process. First, matching scores for all possible target-prediction pairs in a grid cell are stored in a scoring matrix. Then, the best current score in the matrix is used to define a new target-prediction pair that is added to the $\mathds{1}^{\textrm{match}}$ mask and removed from $\mathds{1}^{\textrm{void}}$ if it was not already the case. The full row and column corresponding to the target and detection unit of the best match are masked in the score matrix. This search process is repeated until the score matrix is empty or fully masked. A full example of the association a in a grid cell with the evolution of the score matrix is presented in Fig. 26, associated with Appendix A.5. Interestingly, we observed that our approach presents many similarities with the Kuhn–Munkres algorithm (Kuhn, 1955; Munkres, 1957) that tackles the problem of optimal score-based association, which was not anticipated.
@@
As presented in Sect. 2.2, the best match associations contribute to all subparts of the loss, while the detection units that remained in $\mathds{1}^{\textrm{void}}$ contribute only to the probability and objectness following Eq. 11. The remaining GBNB detection units not being part of either $\mathds{1}^{\textrm{match}}$ or $\mathds{1}^{\textrm{void}}$ do not contribute to the loss. To account for edge cases, size imbalance, or training instability, we added several specific refinements inside our new association process, which we detail in an appendix section dedicated to advanced subtleties of our method (Appendix A.6). The global association function algorithm is presented in Fig. 5 in a way that separates the simplified association process and the advanced association with refinements.